Suppose that as the price of some product increases from $4.00 to $5.00 per unit the quantity supplied rises from 500 to 1000 units per month. The price elasticity of supply for this product is

A)   0.33.
B)   2.0.
C)   2.5.
D)   3.0.
E)   1.0

Suppose an increase in world demand for potash (used in the production of fertilizer) increases the price by 22 percent. Annual Canadian production increases by 33 percent. What is the elasticity of supply of Canadian potash?

A)   0.22
B)   0.33
C)   0.67
D)   1.0
E)   1.5
Answer #1

Answer


Elasticity of supply=(change in quantity/average quantity)/(change in price/average price)
change in quantity=1000-500=500
average quantity=(1000+500)/2=750
change in price=5-4=1
average price=(5+4)/2=4.5
Elasticity of supply=(500/750)/(1/4.5)
=3

The price elasticity of supply for this product is is 3.0
option D
========
Answer


the elasticity of supply of Canadian potash=%change in quantity/%change in price
=33/22
=1.5
the elasticity of supply of Canadian potash is 1.5
Option E